OpenStudy (anonymous):

Y= 2/3x+8 Which are parallel to the one above? A. y=x+6 B. y=-3/2x+11 C. y=4/7x+13 D. Y= 2/3x+13

OpenStudy (anonymous):

Ok, well, there is a way we can tell just from the equation whether two lines are parallel or not. And that by looking at the slope.

OpenStudy (anonymous):

any lines with the same slope are parallel.

OpenStudy (anonymous):

All of these equations are in the y=mx+b format. In this equation m=slope and b=the point where the line intersects the y axis.

OpenStudy (anonymous):

So, can you see which one has the same slope as your equation?

OpenStudy (anonymous):

Sorry guys really stumped here

OpenStudy (anonymous):

That's alright. It's really simple once you get it. Really try to look at the equation y=mx+b. In C for example, m=4/7 and b=13.

OpenStudy (anonymous):

And if m=slope...

OpenStudy (anonymous):

So, in your equation b=8. What would m be?

OpenStudy (anonymous):

Thank you guys just got the response